OpenStudy (anonymous):

Object A travels in the +x-direction before hitting a stationary object B. Afterwards, object A’s x-momentum is 5.7 × 104 kilogram meters/second and its y-momentum is 6.2 × 104 kilogram meters/second. What is the object’s resultant angle of motion with the +x-axis after the collision?

OpenStudy (shamim):

if the motion of object A makes an angle with X axis is\[\theta\]

OpenStudy (shamim):

then the formula will b\[\tan \theta =\frac{ 6.2 \times 10^{4} }{ 5.7 \times 10 ^{4} }\]

OpenStudy (anonymous):

A)42 B)45 C)47 D)48 E)49

OpenStudy (anonymous):

A refrigerator with a weight of 1,127 newtons needs to be moved into a house using a ramp. The length of the ramp is 2.1 meters, and its height is 0.85 meters. What is the percentage efficiency if a force of 496 newtons is applied to the refrigerator?